OpenStudy (anonymous):

What is the y-intercept of y = 0.50(1.06)x?

OpenStudy (mr.math):

Is it \(y=0.50(1.06)x\) or \(y=0.50+1.06x\)?

OpenStudy (mr.math):

Either ways, to find the y-intercept you plug \(x=0\).

OpenStudy (mr.math):

If it's the first you get y=0, which implies the point (0,0).

OpenStudy (aroub):

-I think it's the first one

OpenStudy (mr.math):

He/she had two answers for two questions! @aroub :D

OpenStudy (anonymous):

Its supposed to be y=0.50(1.06)^x

OpenStudy (aroub):

lol non of them! @ Mr.Math

OpenStudy (mr.math):

lol! plug x=0, you get y=0.5. The y-intercept is \((0,0.5)\).

hero (hero):

A ruby

OpenStudy (aroub):

lol you never stop saying it :P I have to tell you something! just a sec :)